Accelerator Update for June 12, 2009 - June 15, 2009 |
The Integrated Luminosity from6/8/09 to 6/15/09 was 50.64 pb-1. Friday June 12 The day shift began with Tevatron (TeV) personnel conducting studies, with Operations monitoring the Recycler stash (302.0E10), with the Antiproton Source stacking (17.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, NuMI, and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store. Main Injector personnel conducted studies. Operations put the TeV through a dry squeeze. Operations put the TeV into HEP shot setup at 12:15 PM. Operations established store 7133 at 1:45 PM with an average initial luminosity (AIL) of 233E30. Operations disabled the NuMI and MTest beam at 9 PM. Operations terminated store 7133 at 9:34 PM. Operations halted the Pbar and Recycler beam at 9:50 PM. Pbar and Recycler personnel began conducting studies. The Pbar lithium lens tripped off at 10:06 PM due to a bad sensor. Pbar was turned off. Operations put the TeV into HEP shot setup and established store 7134 at 11:10 PM with an AIL of 285.9E30. Saturday June 13 The midnight shift began with Operations monitoring store 7134 (225.2E30), and with the Recycler conducting studies. At 12:01 AM, Operations halted beam to all areas in preparation for the shutdown. Recycler personnel continued with their machine studies. Safety system tests began at 7 AM. Recycler studies ended at 12:30 PM. Operators prepared Booster, MI, MI-65, the Pbar rings, Transport, Pre-Vault, Pre-Target, Enclosure C, F-manholes, M01-M05, and MC6 for access. Operations reported finding and isolating a LCW leak on an Accumulator dipole magnet (A2B3). Sunday June 14 The midnight shift began with Operations monitoring store 7134. Operations terminated store 7134 at 12:04 AM. Ops prepared the TeV for access. CHL cryo system technicians started warming up the cryo systems at TeV sectors A3 and D3. Three Linac RF stations (LRF2, 3, and 5) developed problems. Operators contacted the Linac on-call person. ES&H personnel performed surveys everywhere, in preparation for access except for Linac. Locks on the Booster, MI, MI-65, MINOS, and TeV doors were changed over to shutdown cores. The Linac on-call person reported that he had repaired LRF2, but LRF3 continued to have problems. Monday June 15 The midnight shift began with the beam off for shutdown. The Linac on-call person completed his repairs on LRF3, but later LRF2 developed problems again. Tevatron fire alarm testing began around 6 AM. The Plans for Monday and for the rest of the Week Shutdown. |
|
More Information Comments and Suggestions |
| last modified 06/15/2009 email Fermilab |
| Security, Privacy, Legal | |